OpenStudy (anonymous):

find any extrema for sin(x)-cos(x) on (0,2Pi) I already got the second derivative which is -sin(x)cos(x), where do I go from here?

myininaya (myininaya):

what did you get for the first derivative?

OpenStudy (anonymous):

I got F(x)=cos(x)+sin(x)

myininaya (myininaya):

ok did you set that equal to 0 yet to find the critical numbers?

OpenStudy (anonymous):

Well, thats where I am running into a problem. It has been a while since I did calc and after I set it to zero, I don't know how to solve it.

myininaya (myininaya):

so we have cos(x)+sin(x)=0 which means we are looking for when cos(x) and sin(x) have opposite values since cos(x)+sin(x)=0 means cos(x)=-sin(x)

myininaya (myininaya):

recall at pi/4 cos and sin are exactly the same can you tell me an example where they are opposite

OpenStudy (anonymous):

5pi/4?

OpenStudy (anonymous):

I am just guessing here

myininaya (myininaya):

at 5pi/4 they are exactly the same

myininaya (myininaya):

i'm asking you to look at that unit circle and use the definition of opposite

myininaya (myininaya):

i will give you a hint look at the 2nd and 4th quadrants

OpenStudy (anonymous):

Ahh, I see what you mean. 3pi/4?

myininaya (myininaya):

that is one there is one more answer

OpenStudy (anonymous):

and 7pi/4?

myininaya (myininaya):

so the critical numbers are x=3pi/4 , x=7pi/4

myininaya (myininaya):

between 0 and 2pi anyways

OpenStudy (anonymous):

Does it matter any that the problem asks me to use the second derivative?

OpenStudy (anonymous):

second derivative test*

myininaya (myininaya):

also we don't have any extrema values at the endpoints because our endpoints aren't actually included since you have ( ) instead of [ ]

myininaya (myininaya):

you could use the second derivative to determine if they are relative max or relative min or neither

myininaya (myininaya):

f''>0 means it is concave up and you have a relative min example: f=x^2 is concave up and has a relative min and f''=2>0 f''<0 means it is concave down and you have a relative max example f=-x^2 is concave down and has a relative max and f''=-2<0 f''=0 means it is neither relative max or min example f=x^3; f'=3x^2 critical numbers are x=0 f''=6x f''(0)=0 <-- no relative max or min for f=x^3

myininaya (myininaya):

anyways your second derivative you are missing a plus sign somewhere

OpenStudy (anonymous):

and no relative min or max means no extrema, right?

OpenStudy (anonymous):

yea, thats a typo

myininaya (myininaya):

at that particular value yes

OpenStudy (anonymous):

so, to answer my problem, I would solve it with the first derivative, then how would I check it with the second derivative?

OpenStudy (anonymous):

cause my teacher wants me to impliment the second derivative test

myininaya (myininaya):

also local (aka relative) extrema do not occur at endpoints after you find when f'=0 use f'' to determine if you have relative extrema

myininaya (myininaya):

plug in your critical numbers into your second derivative

myininaya (myininaya):

and use that paragraph i gave you above to help you determine what kinda extrema you have

OpenStudy (anonymous):

So, I would have -sin(pi/4) + cos(pi/4) =0 , then do the same for x=7pi/4?

myininaya (myininaya):

err why are you pluggin in pi/4?

OpenStudy (anonymous):

that was one of my critical numbers, right?

myininaya (myininaya):

nope

OpenStudy (anonymous):

"so the critical numbers are x=3pi/4 , x=7pi/4"

myininaya (myininaya):

recall we were looking for when cos(x) and sin(x) were oppsites

myininaya (myininaya):

yes

OpenStudy (anonymous):

So, if they are not the critical numbers, what are they?

myininaya (myininaya):

...

OpenStudy (anonymous):

I'm sorry, I am slow today and I am sure it is testing your patience.

myininaya (myininaya):

i think you need to read above i already told you the critical numbers and you even quoted me on it so i don't understand

OpenStudy (anonymous):

well, I am a little confused, because I thought you were suposed to substitute the critical numbers into the second derivative, right?

myininaya (myininaya):

yep but pi/4 was never a critical number

myininaya (myininaya):

so i asked you why you were plugging in that value

myininaya (myininaya):

pi/4 is not equal to 3pi/4 or 7pi/4

OpenStudy (anonymous):

so, is 7pi/4 the only critical number?

myininaya (myininaya):

why?

myininaya (myininaya):

we said earlier 3pi/4, 7pi/4 are critical numbers

myininaya (myininaya):

they were critical numbers because they made f'=0

OpenStudy (anonymous):

Alright, I think I got it now. Thank you for your help and patience.

myininaya (myininaya):

do you understand why I asked you why you were plugging in pi/4?

OpenStudy (anonymous):

ya, because it wasn't a critical number, right?

myininaya (myininaya):

the only numbers you should be plugging in are the critical numbers which were 3pi/4 , 7pi/4

OpenStudy (anonymous):

Alright, so plug those in and solve them for zero?

myininaya (myininaya):

no solving for zero

myininaya (myininaya):

plug in then use f''>0 means it is concave up and you have a relative min example: f=x^2 is concave up and has a relative min and f''=2>0 f''<0 means it is concave down and you have a relative max example f=-x^2 is concave down and has a relative max and f''=-2<0 f''=0 means it is neither relative max or min example f=x^3; f'=3x^2 critical numbers are x=0 f''=6x f''(0)=0 <-- no relative max or min for f=x^3